A person walks 300m toward due east,then turns back and walks 500m toward due west in a total of 10 minutes. Choose due east to be the positive direction.

a) the total distance covered by the person is:
b) the displacement relative to where the person starts is:
Answer #1

a) Total distance = 300 + 500 = 800 m

b) Relative displacement from starting point = 300 + ( - 500 ) = - 200 m